OpenStudy (anonymous):

Can somebody show me how to do this?? If c carpenters can a build a garage in d days, how many days will it take e carpenters, working at the same rate, to build 2 garages? I know the answer is 2cd/e (thanks to gsmafra ^^), but I would like to see the steps, please. SAT math ftw~

OpenStudy (anonymous):

i think proportions play a part

OpenStudy (anonymous):

Suppose you can make that relation: (Number of carpenters)*(Days of work) = (Garages built)*(Constant) c*d = 1*α (Situation 1) e*n = 2*α (Situation 2) Replacing the constant α = c*d in Situation 2 to find n: e*n = 2*c*d n = 2cd/e

OpenStudy (anonymous):

 e.1.x=c.2.d ====> x= 2cd/e

OpenStudy (anonymous):

this is 3 prportions problem

OpenStudy (anonymous):

is that clear to you?

OpenStudy (anonymous):

Yes. Thanks!

OpenStudy (anonymous):

you are welcome